MATH 541 DIFFERENTIAL TOPOLOGY (3-0)3

Instructor: Sergey Finashin

Content

Manifolds and differential structures. Tangent Space. Vector Bundles. Immersions, submersions, embeddings. Transversality. Sard’s theorem. Whitney embedding theorem. The exponential map and tubular neighborhoods. Manifolds with boundary. Thom’s transversality theorem.

Prerequisite: Consent of the instructor

The first meeting: September 18, Friday, 15-00 (the aim of the meeting is to fix the schedule of lectures, so, please, inform the instructor if you cannot come).